OpenStudy (anonymous):

Sigma n=1 to infinity of cos n*pi/2 ??

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

so: \[\sum_{n=1}^{\infty}\frac{\cos(n \pi)}{2} ?\]

OpenStudy (anonymous):

\[\sum_{1}^{infinity} \cos(nPi/2)\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

What do you want to know? If it converges?

OpenStudy (anonymous):

yes..if it converges and its sum if it converges

OpenStudy (anonymous):

H/o just to make sure I'm rite is it cos(npi/2) or cos(npi)/2? Sorry, I just want to make sure

OpenStudy (anonymous):

first one cos(npi/2)

OpenStudy (anonymous):

> sum((1/2)*cos*n*Pi, n = 1 .. infinity); "(->)" infinity = 0

OpenStudy (anonymous):

@inik: doesn't it depend on n??

OpenStudy (anonymous):

Actually, it does not converge.

OpenStudy (anonymous):

yes , i second male violence

OpenStudy (anonymous):

It doesn't converge. Because if you take the lim n->infinity you get D.N.E. For a sum to converge, the lim n->infinity has to be ZERO.

OpenStudy (anonymous):

if n is odd even doesn't it converge?

OpenStudy (anonymous):

If n is odd it does. To zero. But if it is even it switches between 1 and -1.

OpenStudy (anonymous):

so what must be the final answer..infinity and it diverges?

OpenStudy (anonymous):

Not necessarily infinity. But it just diverges. You can diverge without being infinite.

OpenStudy (anonymous):

Thanks :)

OpenStudy (anonymous):

No problem :P

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}(\cos n \Pi)/2^{n}\]

OpenStudy (anonymous):

converges or diverges? if converges what is the sum?

OpenStudy (anonymous):

Okay. So, we agree, that for no matter what value of n you plug in. cos(npi) is either +1 or -1? Starting with n=1 you get cos(pi)=-1 so: \[\sum_{n=1}^{\infty}\frac{\cos(n \pi)}{2^n}=\sum_{n=1}^{\infty}\frac{(-1)^n}{2^n}\]. Well, that is the same as: \[\sum_{n=1}^{\infty}(\frac{-1}{2})^{n}\]. This is ALMOST a geometric series. You notice if you do n=1 the series DOES NOT start at zero. So reindex it: \[\sum_{n=1}^{\infty}(\frac{-1}{2})(\frac{-1}{2})^{n-1}\]. So you know a geometric series converges if the absolute value of the radius is less than 1 so: \[\sum_{n=1}^{\infty}ar^{n-1}=\frac{a}{1-r} \iff |r|<1\]. Since r=(-1/2)=>|-1/2|<1. The series converges to: \[\frac{\frac{-1}{2}}{1-(\frac{-1}{2})}\]. or -1/3

OpenStudy (anonymous):

Sorry that took me so long to type. Latex can be a feather sometimes xP

OpenStudy (anonymous):

Make sense? :O

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

OMG! got it now! Thanks a lot!

OpenStudy (anonymous):

No problem :) Just post any questions you have Calculus is my thing xP

OpenStudy (anonymous):

maleviolence19 dat was a stud proof !

OpenStudy (anonymous):

Haha, Thank you! I try :D

OpenStudy (anonymous):

Is sigma n=1 to infinity of (3^n -1) divided by 2^n divergent? just tell me if it is right or wrong!

OpenStudy (anonymous):

yeap!

OpenStudy (anonymous):

It is divergent. Because you can break it up: \[\sum_{n=1}^{\infty}\frac{3^n-1}{2^n}=\frac{3}{2}\sum_{n=1}^{\infty}(\frac{3}{2})^{n-1}-\frac{1}{2}\sum_{n=1}^{\infty}(\frac{1}{2})^{n-1}\]. You know that the left one diverges and the right one converges. But a divergent added to a convergent still diverges!!!

OpenStudy (anonymous):

ty!

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

You could also do the ratio test :P

OpenStudy (anonymous):

You are so smart! are you a calc prof or sth?

OpenStudy (anonymous):

Nah, haha. I am a math major senior in college though :P

OpenStudy (anonymous):

Nice!

OpenStudy (anonymous):

Yeah :D

OpenStudy (anonymous):

Sigma n=1 to infinity (8^n)/(8^n)..diverges right